# Understanding Differences in News Article Interaction Patterns on Facebook: Public vs. Private Sharing with Varying Bias and Reliability

### v2

@@
Abstract
The rapid growth of news dissemination and user engagement on social media has raised concerns about the influence and societal impact of biased and unreliable information. As a response to these concerns, a substantial body of research has been dedicated to understanding how users interact with different news. However, this research has primarily analyzed publicly shared posts. With a significant portion of engagement taking place within Facebook’s private sphere, it is therefore important to also consider the private posts. In this paper, we present the first comprehensive comparison of the interaction patterns and depth of engagement between public and private posts of different types of news content shared on Facebook. To compare these patterns, we gathered and analyzed two complementary datasets: the first includes interaction data for all Facebook posts (private + public) referencing a manually labeled collection of over 19K news articles, while the second contains only interaction data for public posts tracked by CrowdTangle. As part of our methodology, we introduce several carefully designed data processing steps that address some critical aspects missed by prior works but that (through our iterative discussions and feedback with the CrowdTangle team) emerged as important to ensure fairness for this type of study. Our findings highlight significant disparities in interaction patterns across various news classes and spheres. For example, our statistical analysis demonstrates that users engage significantly more deeply with news in the private sphere compared to the public one, underscoring the pivotal role of considering both the public and private spheres of Facebook in future research. Beyond its scholarly impact, the findings of this study can benefit a wide range of stakeholders, including Facebook content moderators, regulators, and policymakers, contributing to a healthier online discourse.
keywords:
News sharing, Facebook, User engagement, Bias, Reliability, Public vs private
1 Introduction
Social media has become a dominant medium for news consumption and sharing, with Facebook being the most popular and widely used platform for these purposes [1, 2]. For example, according to a 2021 Pew Research Center survey [2], 66% of all Americans regularly use Facebook, and almost half of them (31% of the US population) regularly get their news via the platform. Other popular platforms that users regularly use for news consumption include YouTube (22% of the US population), Twitter (13%), and Instagram (11%). This significant reach highlights the potential impact that news consumption and sharing on Facebook (and social media in general) can have on the opinions and thoughts of both individuals and the society [3].
@@ 1 Introduction
Unfortunately, not all news are reliable. With an increasing amount of misinformation being circulated on Facebook, it is crucial to understand how users engage with news of varying reliability [4]. Improved insights and understanding into these dynamics are expected to benefit media researchers, journalists, content moderators, and policymakers, which choices based on these insights, in turn, are expected to influence public opinion and the behavior of regular users [5, 6].
@@
Previous research has extensively examined public user engagement dynamics with different news content and the factors affecting this engagement  [7, 8, 9, 10, 11, 12]. However, with a growing privacy inclination of Facebook users and a rise in user engagement in private spaces of Facebook [13, 14], it is becoming increasingly important to understand these dynamics also with regards to private sharing. Yet, the current literature predominantly focuses on public posts, leaving gaps in our understanding of engagement differences between these realms. Most importantly, prior research has not studied the variations in public vs. non-public interaction with regards to (1) reliable and unreliable news content, (2) news with different political biases, or (3) the depth of the interactions associated with each of these types of posts and content types. These are the primary gaps addressed in this paper.
@@
In this paper, we examine the differences in public and private sharing of news articles on Facebook, with an emphasis on users’ interactions with these articles. Specifically, we investigate how news articles written with different bias and reliability are shared and engaged with, as well as the depth of these interactions. In this context, bias refers to a tendency for news articles to exhibit partiality or favoritism towards particular groups or ideas (e.g., left or right on the political spectrum), while reliability refers to the accuracy and credibility of the information presented [15] (e.g., fake or true). While most previous research focuses on either bias or reliability, we consider both dimensions, as they have been found related, but yet each has its own ability to influence news article sharing behaviors and affect the quality and diversity of information that users encounter on social media [16, 17]. On the other hand, by considering whether an interaction is deep or shallow, we can identify whether there are areas (private or public) where users engage more deeply with specific content types (e.g., fake news). This is important since the depth of user engagement can significantly influence the success of an article. For example, an article may have greater success if people share and comment on the posting of the article than if they only press Facebook’s “like” button [1].
Research questions and methodology: With the above goals in mind, we designed our study to address the following research questions:
@@ 2. RQ2
We took several steps to address these previously unaddressed questions in as controlled matter as possible. First, in contrast to most prior works that use publisher-level labeling of news (e.g., Adelson et al. [7]; Horne et al. [18]), we employ article-level labeling of bias and reliability. By doing so, we capture that not all articles by a publisher have the same bias or reliability, recognizing that the publisher (source) is only one factor [19] in the bias/reliability. In particular, we consider the bias and reliability of articles ($N=19,050$), meticulously selected from a substantial pool of publishers ($N=1,121$) and 30K+ manually labeled news articles. Here, the original labeling was provided by Ad Fontes Media [20], which rates both news sources and individual articles based on their level of bias and reliability. Second, through active back-and-forth discussions and feedback with the Crowdtangle team, we developed and implemented a methodology using their Chrome addon that allows us to obtain simultaneous interaction statistics for a representative sample of publicly shared Facebook posts linking these articles, as well as across all Facebook posts (including both public and private posts) linking these articles. Finally, using this unique dataset, we perform a comparison of interaction dynamics when users share news articles with different levels of bias and reliability, both publicly and privately.
@@
Example beneficiaries: Our methodology and findings contribute to a better understanding of the dynamics of news sharing and interaction on Facebook and provide valuable insights for multiple stakeholders. For example, considering the significant impact that social media interactions have on content and platform popularity [21], our findings on how bias and reliability influence user engagement in public vs. private spheres can aid media researchers and journalists. These insights can inform more effective news content creation and distribution strategies for enhanced audience engagement. For content moderators of platforms such as Facebook, the analysis offers data-driven guidance on where moderation efforts may be most needed to curb the spread of problematic content. For policymakers and regulators, our findings illustrate how user privacy settings impact news content engagement types and interaction depth. Notably, highly unreliable news receives more private discussion engagement share compared to the broader unreliable category. These results underscore how platform design choices can influence selective exposure and engagement, potentially exacerbating issues like ideological polarization and misinformation spread.
Roadmap: Sect. 2 explains our multi-faceted research design, including news article selection, labeling, interactions data collection, and the processing steps taken to ensure fair comparisons between public and private engagement. The resulting dataset is summarized in Sect. 3. Sect. 4 analyzes public vs. private dynamics both from a high-level aggregated perspective (Sect. 4.1) and then using a detailed statistical analysis (Sect. 4.2). In Sect. 5, we turn our attention to differences in the depth of interactions. Sect. 6 concludes our analysis with a correlation-based analysis between public and private interactions. Discussion of the key findings is provided in Sect. 7. Finally, related works (Sect. 8) and conclusions (Sect. 9) are presented.
2 Research Design, Data Collection and Limitations
@@
  • 1.
    Article selection and labeling: We first obtain the URLs to a substantial number of news articles, which have been classified by Ad Fontes Media [20] with regards to both their individual bias and reliability. We then decide category thresholds and label each article into one of three bias classes and one of four reliability classes.
  • 2.
    Collection of interactive data: After careful preprocessing of the URLs, we use the CrowdTangle browser extension to obtain interaction data for two sets of posts linked to these URLs (1) public and (2) combined (public + private), where the first set is a subset of the second. To address some limitations of the API and to ensure fair comparison of the sets, we leave a four-month gap between the latest labeled article (Nov. 1, 2022) and the primary data collection (Mar. 2023), as well as apply some additional post-processing (e.g., examining the actual posts), filtering (e.g., based on thresholds determined via communication with the CrowdTangle team), and collect some complementing data directly from CrowdTangle (e.g., to address limitations of the extension API).
  • 3.
    Data analysis: Finally, we use the final dataset (capturing the two sets) to compare the properties of the public vs. combined (private + public) sets and discover statistically significant differences between the interaction patterns of public and private posts. Here, it should also be noted that the combined set is dominated by private posts.
We next provide detailed descriptions of the initial two steps, before presenting our analysis results (step 3) in the subsequent sections.
2.1 News Article Selection and Labeling
Selection of articles: There are several independent initiatives that assess the bias and/or reliability of individual news articles and/or news sources. Examples of such evaluation efforts include Media Bias Fact Check [22], Ad Fontes Media, AllSides [23], and NewsGuard [24]. Among these, we opted to use data from Ad Fontes Media for the following main reasons: (1) it has been widely used in previous research [25, 26, 27, 12], (2) they evaluate individual news articles (not only the news publishers), (3) each assessed article receives a score for both bias and reliability, (4) they offer a transparent evaluation methodology, which is published and explained in a white paper [28], and finally (5) the dataset comprises a large set of news contents (including news articles, podcasts and TV shows) randomly selected from a diverse set of publishers.
We obtained all evaluated news from Ad Fontes Media as of Nov. 1, 2022. After filtering the podcasts and TV shows, this set included 31,446 news article URLs from 1,121 publishers. We also filtered a few URLs that (1) primarily served as event reporting pages (e.g., https://www.nola.com/news/hurricane) and (2) articles for which Ad Fontes Media had changed their rating (identified through complementary documents published by Ad Fontes Media). After excluding these articles, the dataset comprised 31,408 articles.
@@ 2.1 News Article Selection and Labeling
Figure 1: CDF of bias scores
Figure 2: CDF of reliability scores
@@
For our analysis, we categorized the bias and reliability scores into distinct classes. While previous work mainly focuses on binary classification of news (e.g., Left vs. Right or Fake vs. True), we opted for a more granular approach by defining three classes of bias (Left, Right, and Center) and four classes of reliability (Most-unreliable, Unreliable, Reliable, and Most-reliable). Figures 1 and 2 show the Cumulative Distribution Function (CDF) of the bias and reliability scores of the articles included in our analysis (after additional filtering steps, explained later in this section) together with the threshold values and class labels used to define each class.
@@
With our labeling, anything with a bias score below -6 is called Left-biased, and anything above 6 is called Right-biased, with the range (-6,6) capturing the Center class. We note that Ad Fontes Media also refers to this range as “Middle or Balanced Bias” [20]. As observed, this class accounts for most samples in our dataset, representing 58% (11,094 out of 19,050 articles in the final dataset). With this split, the two biased classes have similar, close to 20%, shares.
@@
For the reliability classes, we assigned the Most-unreliable class to cover the lowest range (0, 32). According to Ad Fontes Media’s terminology [20], these values typically correspond to articles “which contain inaccurate and misleading info, selective stories or have opinion and wide variation in their reliability”. Similarly, the range (48-64), chosen for the Most-reliable class, encompasses articles characterized by “thorough and original fact reporting” due to Ad Fontes terminology. Finally, the threshold of 40, between Reliable and Unreliable, matches the mid-point between these two classes as well as where Ad Fontes make their split between “wide variation in reliability” and “mix of fact reporting and analysis”.
2.2 Collection of Interaction Data
@@
Preprocessing of URLs: Before collecting the interaction data associated with each article (URL), we (1) replaced the link shorteners with their corresponding URLs and (2) converted all article URLs to their canonical forms. During the conversion process, we primarily removed all unnecessary URL parameters (those sent after the ”?” symbol) except for the ones essential to the canonical form (e.g., “id” parameters are sometimes part of the canonical form). As another step in identifying the canonical forms, we replaced all the web archive links in the Ad Fontes dataset with the original links.
@@
Primary data collection for the two sets of posts: We used the CrowdTangle browser extension [29] to collect the interaction data for the Facebook posts sharing each news article of interest. In addition to interaction statistics for the public posts that CrowdTangle (owned by Facebook) indexes and tracks over time (their database, available for and used by researchers, contains longitudinal interaction data for indexed posts), this extension provides access to the life-time interaction statistics, obtained thus far, across all posts (private + public). In this paper, we use this extension to collect statistics about all interactions associated with these two sets of posts: combined and public, where the combined captures all interactions of all posts (public + private) and the public is a subset of the combined (only capturing public posts).
@@
To collect the data for each URL, we used the “Download” option within the extension’s interface when browsing each article. This allowed us to easily download a CSV file containing the interactions data for both sets. As expected from a privacy standpoint, only the aggregated statistics are provided for the combined set. In contrast, every post and their individual interaction data are provided for the public set.
@@
At a high level, the combined category aggregates the interactions (e.g., likes, shares, comments, etc.) of all Facebook posts referencing the article’s URL, regardless of the post’s privacy settings. This includes interactions on posts with limited privacy settings, such as “Friends only” or “Only Me”. In contrast, the public category, includes all interactions with the public posts tracked by CrowdTangle, which, due to heavy-tailed characteristics and significant coverage, captures most interactions with public Facebook posts. As supporting examples, all posts of all US-based public groups with 2K+ members are indexed, and in 2021 over 99% of all Facebook pages with at least 25K likes were indexed [30]. While some public posts are not indexed, we note that their indexing does not consider any bias/reliability criteria, is expected to capture the majority of public posts, and the majority of the posts in the combined set are private.
@@
Timeline and limitations: We collected the interaction data in Mar. 2023, ensuring at least a four-month gap between the publication date of the newest article in our dataset and our interaction data collection. This length of interval was chosen to ensure that we capture most of the posts and interactions for the URLs in our dataset.
@@
As mentioned in [31] and confirmed via our communication with CrowdTangle, the information from the CrowdTangle addon originates from two different sources: (1) public interaction data is obtained from the CrowdTangle database, while (2) combined interaction data is derived from Facebook’s Graph API. Consequently, variations in data capturing times might cause some URLs to exhibit higher public than combined interactions [31]. Furthermore, as discussed in [31], Facebook’s aggregation data restarts the interaction count whenever websites update their connection schemes (e.g., switching from HTTP to HTTPS) after publishing a URL, while CrowdTangle continues accumulating data. Although the discrepancy is not substantial (9% in our initial dataset), we aimed to mitigate the effect of these differences where possible (described next).
Further filtering and data processing for fairer comparisons: After consulting with the CrowdTangle team regarding all numerical discrepancies that we observed, we identified two additional reasons for the discrepancies. First, as stated in [32], for privacy reasons, for the combined data (that include private posts) “the values are intentionally not precise, but you can be confident they accurately reflect user engagement with a URL.” Discussing the identified discrepancy for such cases, the CrowdTangle team confirmed that combined interaction data with counts below 100 is not reliable; however, that data with counts above 100 provides a reliable estimate of public + private interactions. For this reason, we removed all URLs in our dataset with combined interactions below 100. Although this reduced our dataset to 19,505 articles, it is important to note that this group of URLs accounted for less than 4% of public interactions in our dataset. Furthermore, we note that the public interaction data for this group still remains reliable.
Second, the CrowdTangle team informed us that the Facebook Graph API only considers posts with the URL attached to the post (i.e., those displaying a preview of the URL) when counting interaction with the URL, while CrowdTangle’s algorithm includes any post containing a linked URL, regardless of attachment status. Taking this into account, in our final step toward enhancing interaction data quality, we chose not to rely on the aggregated data from the CrowdTangle extension (which sums the interactions of all posts mentioned in the extension). Instead, we examined each post to determine whether the related URL was attached or not, and if so, we included them in the sum of public interactions for that URL. [^1] The attachment status of all the posts and URLs will be published with the camera-ready version for the interested researchers. This resulted in a more accurate comparison between combined and public interactions. After these processing steps, we had only 455 URLs for which public interactions exceeded combined interactions, which were excluded from our final dataset. The primary causes of these discrepancies were beyond the scope of our research to address (e.g., changes in URL, publisher protocol schemes, or different data capturing times as mentioned above).
While some of the pre- and post-processing steps described above require significant effort, they help ensure an accurate and fair comparison between the public and combined sets in such a way that we can provide conclusive insights into the relative sharing patterns of public vs. private posts.
Enhancing the data with additional posts: Another limitation of the CrowdTangle extension is that it only allows us to retrieve data for up to 500 posts per URL. To address this limitation, for the URLs with more than 500 posts in the public interaction category, we obtained additional data directly from CrowdTangle. However, these URLs constitute only 42 instances in our dataset and will have no significant impact on the reproducibility of our statistical analysis.
3 Dataset
High-level summary: After applying the aforementioned filtering steps, we have in total 19,050 articles from 1,121 news outlets remaining in our dataset. These articles have been shared in 253,350 posts, which combined are responsible for 225,282,745 interactions (out of which 70,428,316 are public). Table 1 summarizes the dataset.
Table 1: Summary of the final dataset
Type Articles Publishers Posts Combined interactions Public interactions
Number 19,050 1,121 253,350 225,282,745 70,428,316
Table 2: Statistics based on Bias and Reliability

### v3

# Public versus Less-Public News Engagement on Facebook: Patterns Across Bias and Reliability

@@
Abstract.
The rapid growth of social media as a news platform has raised significant concerns about the influence and societal impact of biased and unreliable news on these platforms. While much research has explored user engagement with news on platforms like Facebook, most studies have focused on publicly shared posts. This focus leaves an important question unanswered: how representative is the public sphere of Facebook’s entire ecosystem? Specifically, how much of the interactions occur in less-public spaces, and do public engagement patterns for different news classes (e.g., reliable vs. unreliable) generalize to the broader Facebook ecosystem?
This paper presents the first comprehensive comparison of interaction patterns between Facebook’s more public sphere (referred to as public in paper) and the less public sphere (referred to as private). For the analysis, we first collect two complementary datasets: (1) aggregated interaction data for all Facebook posts (public + private) for 19,050 manually labeled news articles (225.3M user interactions), and (2) a subset containing only interactions with public posts (70.4M interactions). Then, through discussions and iterative feedback from the CrowdTangle team, we develop a robust method for fair comparison between these datasets.
Our analysis reveals that only 31% of news interactions occur in the public sphere, with significant variations across news classes. Engagement patterns in less-public spaces often differ, with users, for example, engaging more deeply in private contexts. These findings highlight the need to examine both public and less-public engagement to fully understand news dissemination on Facebook. The observed differences hold important implications on content moderation, platform governance, and policymaking, contributing to healthier online discourse.
journalyear: 2025copyright: ccconference: Proceedings of the 17th ACM Web Science Conference 2025; May 20–24, 2025; New Brunswick, NJ, USAbooktitle: Proceedings of the 17th ACM Web Science Conference 2025 (Websci ’25), May 20–24, 2025, New Brunswick, NJ, USAdoi: 10.1145/3717867.3717912isbn: 979-8-4007-1483-2/25/05ccs: Information systems World Wide Web
1. Introduction
Social media has emerged as an increasingly important news source, with Facebook maintaining a prominent position among social platforms  (Newman et al., 2023; Christopher Aubin, 2024). For example, according to a 2024 Pew Research Center report (Christopher Aubin, 2024), 33% of U.S. adults regularly get news on Facebook, similar to YouTube (32%), and notably ahead of Instagram (20%) and Twitter (12%). This extensive reach highlights the potential impact of news sharing on Facebook in shaping both individual opinions and broader societal discourse (Thompson et al., 2020).
@@ 1. Introduction
Unfortunately, not all news are reliable. With an increasing amount of misinformation being circulated on Facebook, it is crucial to understand how users engage with news of varying reliability (Geeng et al., 2020). Understanding of these dynamics are expected to benefit media researchers, journalists, content moderators, and policymakers, whose choices based on these insights, in turn, are expected to influence public opinion and the behavior of regular users (Nah and Yamamoto, 2018; Jones et al., 2017).
@@
Previous research has extensively examined public user engagement dynamics with different news content and the factors affecting this engagement  (Edelson et al., 2021; Aldous et al., 2019; Brena et al., 2019; Marwick, 2018; Vitak et al., 2011). However, with a growing privacy inclination of Facebook users and a rise in user engagement in private spaces of Facebook (Dey et al., 2012; Lottridge and Bentley, 2018), it is becoming increasingly important to understand these dynamics also with regards to private sharing. Yet, the current literature is centered around more public posts, leaving gaps in our understanding of engagement differences between these realms. Most importantly, prior research has not studied the dynamics of public vs. private interaction.
@@
To address this gap, in this paper we present the first comprehensive comparison of the news article sharing and user interaction patterns seen on Facebook’s more public sphere (referred to as public in paper) versus in the less public sphere (for simplicity referred to as private). Specifically, we investigate differences in how news articles written with different bias and reliability are shared and engaged with, as well as the depth of these interactions. In this context, bias refers to a tendency for news articles to exhibit partiality or favoritism towards particular groups or ideas (e.g., left or right on the political spectrum), while reliability refers to the accuracy and credibility of the information presented (e.g., fake or true). While most previous research on news engagement has primarily focused on either bias or reliability (and on public engagement), we consider both dimensions, as they have been found related, but yet each has its own ability to influence news article sharing behaviors and affect the quality and diversity of information that users encounter on social media (Guess et al., 2018; Pennycook et al., 2018).
Furthermore, by analyzing whether interactions are deep or shallow, we can determine if users engage more deeply with certain content types (e.g., fake news) in private or public spaces. This distinction is crucial, as deeper engagement—such as sharing or commenting—can significantly enhance an article’s impact compared to shallow actions like pressing Facebook’s “like” button (Newman et al., 2023).
Research questions and methodology: Guided by the goal of better understanding how representative the public sphere of Facebook’s ecosystem (typically studied in prior works) is of the private sphere and the ecosystem as a whole, we designed our study to address the following research questions:
@@ 2. RQ2
We took several steps to address these previously unaddressed questions in as controlled matter as possible. First, in contrast to most prior works that use publisher-level labeling (e.g., Adelson et al. (Edelson et al., 2021); Horne et al. (Horne et al., 2019)), we employ article-level labeling of bias and reliability. By doing so, we capture that not all articles by a publisher have the same bias or reliability, recognizing that the publisher (source) is only one factor (Pidikiti et al., 2020) in the bias/reliability. In particular, we consider the bias and reliability of articles ($N=19,050$), meticulously selected from a substantial pool of publishers ($N=1,121$) and 30K+ manually labeled news articles. Here, the original labeling was provided by Ad Fontes Media (Ad Fontes Media, 2024), which provides bias and reliability of news articles. Second, through active back-and-forth discussions and feedback with the Crowdtangle team, we developed and implemented a methodology using their Chrome addon that allows us to obtain simultaneous interaction statistics for a representative sample of publicly shared Facebook posts linking these articles, as well as across all Facebook posts (including both public and private posts) linking these articles. Finally, using this unique dataset, we perform a comparison of interaction dynamics when users share news articles with different levels of bias and reliability, both publicly and privately.
@@
Example beneficiaries: Our methodology and findings contribute to a better understanding of news sharing and interaction dynamics on Facebook, offering valuable insights for various stakeholders. Media researchers and journalists can use our analysis of how bias and reliability shape engagement in public and private spheres for improved content creation and distribution strategies. For content moderators, the analysis offers data-driven guidance on prioritizing efforts to curb the spread of problematic content. For policymakers, we highlight how privacy settings influence engagement patterns, with highly unreliable news, for example, garnering more engagement in the private sphere. These results underscore how platform design choices can influence selective exposure and engagement, potentially exacerbating issues like ideological polarization and misinformation spread.
Roadmap: Sect. 2 explains our research design, including news article selection, labeling, data collection, and the processing steps taken to ensure fair comparisons. The resulting dataset is summarized in Sect. 3. Sect. 4 analyzes public vs. private dynamics both from a high-level aggregated perspective (Sect. 4.1) and then using a detailed statistical analysis (Sect. 4.2). In Sect. 5, we turn our attention to differences in the depth of interactions. Topic and top-publishers analysis are provided in Sects. 6 and 7. Finally, related works (Sect. 8) and conclusions (Sect. 9) are presented.
2. Research Design, Data Collection and Limitations
@@
Part 1. Article selection and labeling: We first obtained bias and reliability scores for the articles evaluated by Ad Fontes Media (Ad Fontes Media, 2024). Subsequently, each article was categorized into one of three bias classes and one of four reliability classes based on these scores.
Part 2. Collection of interactive data: After careful preprocessing of the URLs, we used the CrowdTangle browser extension to obtain interaction data for two sets of posts linked to these URLs (1) public and (2) combined (public + private). To address some limitations of the API and to ensure fair comparison of the sets, we leave a four-month gap between the latest labeled article (Nov. 1, 2022) and the primary data collection (Mar. 2023), as well as apply some additional post-processing (e.g., examining the actual posts), filtering (based on thresholds determined via discussions with the CrowdTangle team), and collect some complementing data directly from CrowdTangle (to address limitations of the extension API).
Part 3. Data analysis: Finally, we use the final dataset (capturing the two sets) to compare the properties of the public vs. combined (private + public) sets and identify significant statistical differences between the interaction patterns of public and private posts. We next provide details of the initial two steps, before presenting our analysis results (Part 3) in the subsequent sections.
2.1. News Article Selection and Labeling
Selection of articles: There are several independent initiatives that assess the bias and/or reliability of individual news articles and/or news sources. Examples of such evaluation efforts include Media Bias Fact Check (Media Bias Fact Check, 2024), Ad Fontes Media, AllSides (AllSides, 2023), and NewsGuard (NewsGuard, 2024). Among these, we opted to use data from Ad Fontes Media for the following main reasons: (1) it has been widely used in previous research (Haq et al., 2022; Huszár et al., 2022; Knutson et al., 2024), (2) they evaluate individual news articles (not only the news publishers), (3) each assessed article receives a score for both bias and reliability, (4) they offer a transparent evaluation methodology, which is published and explained in a white paper (Otero, 2021), and finally (5) the dataset comprises a large set of news contents randomly selected from a diverse set of publishers.
We obtained all evaluated news from Ad Fontes Media as of Nov. 1, 2022. This included 31,446 news article URLs from 1,121 publishers. After filtering out URLs serving as event reporting pages (e.g., https://www.nola.com/news/hurricane) and articles with updated ratings (per Ad Fontes Media’s documentation), the final dataset included 31,408 articles.
@@ 2.1. News Article Selection and Labeling
For our analysis, we categorized the bias and reliability scores into distinct classes. While previous work mainly focuses on binary classification of news (e.g., Left vs. Right or Fake vs. True), we opted for a more granular approach by defining three classes of bias (Left, Right, and Center) and four classes of reliability (Most-unreliable, Unreliable, Reliable, and Most-reliable). Figures 1(a) and 1(b) show the Cumulative Distribution Function (CDF) of the bias and reliability scores of the articles included in our analysis (after additional filtering steps, explained later in this section) together with the threshold values and class labels used to define each class.
@@
With our labeling, anything with a bias score below -6 is called Left-biased, and anything above 6 is called Right-biased, with the range (-6,6) capturing the Center class. We note that Ad Fontes Media also refers to this range as “Middle or Balanced Bias” (Ad Fontes Media, 2024). As observed, this class represents most samples in our dataset, representing 58% (11,094 out of 19,050 articles in the final dataset). With this split, the two biased classes have similar, close to 20%, shares.
@@
For the reliability classes, we assigned the Most-unreliable class to cover the lowest range (0, 32). According to Ad Fontes Media’s terminology (Ad Fontes Media, 2024), these values typically correspond to articles “which contain inaccurate and misleading info, selective stories or have opinion and wide variation in their reliability”. Similarly, the range (48-64), chosen for the Most-reliable class, encompasses articles characterized by “thorough and original fact reporting” due to Ad Fontes terminology. Finally, the threshold of 40, between Reliable and Unreliable, matches the mid-point between these two classes as well as where Ad Fontes make their split between “wide variation in reliability” and “mix of fact reporting and analysis”.
2.2. Collection of Interaction Data
@@
Preprocessing of URLs: Before collecting the interaction data associated with each article (URL), we (1) expanded link shorteners and (2) converted URLs to canonical forms. During the conversion process, we primarily removed unnecessary URL parameters after “?” except those essential to canonical forms (e.g., “id” parameters are sometimes part of the canonical form). We also replaced web archive links in the Ad Fontes Media’s dataset with original links.
@@
Primary data collection for the two sets of posts: Prior to its shutdown in Aug. 2024, CrowdTangle (owned by Facebook) provided access to interaction statistics for public posts, which it indexed in a database widely used by researchers. Additionally, their Chrome extension offered access to the life-time interaction statistics available at the time, encompassing all posts (public + private). In this study, we utilized the extension to collect statistics for two datasets: combined and public, where the combined dataset includes all interactions from all posts (public + private), and the public dataset represents a subset of the combined dataset, containing only interactions with publicly accessible posts.
(a) Bias scores
(b) Reliability scores
Figure 1. CDFs of scores for (a) bias and (b) reliability.
@@
To collect the data for each URL, we used the “Download” option within the extension’s interface when browsing each article. The downloaded CSV file contained the interaction data for both sets. As expected from a privacy standpoint, only the aggregated statistics are provided for the combined set. In contrast, every post and its individual interaction data are provided for the public set.
@@
At a high level, the combined category aggregates the interactions (e.g., likes, shares, comments, etc.) of all Facebook posts referencing the article’s URL, regardless of the post’s privacy settings. This includes interactions on posts with limited privacy settings, such as “Friends only” or “Only Me”. In contrast, the public category, includes all interactions with the public posts tracked by CrowdTangle, which, due to heavy-tailed characteristics and significant coverage, captures most interactions with public Facebook posts. As supporting examples, all posts of all US-based public groups with 2K+ members are indexed, and in 2021 over 99% of all Facebook pages with at least 25K likes were indexed (Tess, 2021).
@@
Timeline and limitations: We collected the interaction data in Mar. 2023, ensuring at least a four-month gap between the publication date of the newest article in our dataset and our interaction data collection. This interval ensured capturing most posts and interactions for the URLs. This four-month window is notably conservative, as research has shown that most interactions occur within days of publication (Thorgren et al., 2024; Pfeffer et al., 2023). Although this conservative window may have overlooked interactions with content that was later removed or deleted, we are unable to assess the extent of this effect—if it exists—due to CrowdTangle’s shutdown in Aug. 2024.
@@
As mentioned in (Garmur, 2022), the information from the CrowdTangle addon originated from two different sources: (1) public interaction data is obtained from the CrowdTangle database, while (2) combined interaction data is derived from Facebook’s Graph API. Consequently, variations in data capturing times might cause some URLs to exhibit higher public than combined interactions.
Furthermore, as confirmed by the CrowdTangle team, Facebook’s aggregation data restarts the interaction count whenever websites update their connection schemes (e.g., switching from HTTP to HTTPS) after publishing a URL, while CrowdTangle continued accumulating data. Although the discrepancy is not substantial (9% in our initial dataset), we aimed to mitigate the effect of these differences where possible (described next).
Table 1. Statistics based on Bias and Reliability.
@@
Total 19,050 225,282,745 70,428,316
Figure 2. Interactions per article
Further filtering and data processing for fairer comparisons: After consulting with the CrowdTangle team regarding all numerical discrepancies that we observed, we identified two additional reasons for the discrepancies. First, as stated in (Meta, 2023), for privacy reasons, for the combined data (that include private posts) “the values are intentionally not precise, but you can be confident they accurately reflect user engagement with a URL.” Discussing the identified discrepancy for such cases, the CrowdTangle team confirmed that combined interaction data with counts below 100 is not reliable; however, that data with counts above 100 provides a reliable estimate of public + private interactions. For this reason, we removed all URLs in our dataset with combined interactions below 100. Although this reduced our dataset to 19,505 articles, it is important to note that this group of URLs accounted for less than 4% of public interactions in our dataset. Furthermore, we note that the public interaction data for this group still remains reliable.
Second, the CrowdTangle team informed us that the Facebook Graph API only considers posts with the URL attached to the post (i.e., those displaying a preview of the URL) when counting interaction with the URL, while CrowdTangle’s algorithm includes any post containing a linked URL, regardless of attachment status. Taking this into account, in our final step toward enhancing interaction data quality, we chose not to rely on the aggregated data from the CrowdTangle extension (which sums up the interactions of all posts mentioned in the extension). Instead, we examined each post to determine whether the related URL was attached or not, and if so, we included them in the sum of public interactions for that URL.
This resulted in a more accurate comparison between combined and public interactions. After these processing steps, we had only 455 URLs for which public interactions exceeded combined interactions, which were excluded from our final dataset. The primary causes of these discrepancies were beyond the scope of our research to address (e.g., changes in URL, publisher protocol schemes, or different data capturing times as mentioned above).
Enhancing the data with additional posts: Another limitation of the CrowdTangle extension was its restriction to retrieving data for up to 500 posts per URL. To address this, we retrieved additional data from CrowdTangle for URLs with over 500 posts in the public interaction category. While some of the pre- and post-processing steps described above require significant effort, they help ensure an accurate and fair comparison between the public and combined sets in such a way that we can provide conclusive insights into the relative sharing patterns of public vs. private posts.
3. Dataset
High-level summary: After applying the aforementioned filtering steps, we have in total 19,050 articles from 1,121 news outlets remaining in our dataset. These articles have been shared in 253,350 posts, which combined are responsible for 225,282,745 interactions (out of which 70,428,316 are public).